# Trust Me, I Can Convince You: The Contextualized Argument Appraisal Framework

### v1

@@ Abstract
Emotions, which influence how convincing an argument is, are developed in context of the self and sender, and therefore require modeling the cognitive evaluation process. While binary emotionality has been studied in argument mining, and the cognitive appraisal has been modeled in general emotion analysis, these fields have not been brought together yet. We therefore propose the Contextualized Argument Appraisal Framework that contextualizes the interplay between the sender, receiver, and argument. It includes emotion labels, appraisals, such as argument familiarity, response urgency, and expected effort, as well as convincingness variables. To evaluate the framework and pave the way to computational modeling, we perform a study in a role-playing scenario, mimicking real-world exposure to arguments, asking participants to disclose their emotion, explain the main cause, the argument appraisal, and the perceived convincingness. To consider the subjective nature of such annotations, we also collect demographic data and personality traits of both the participants and the perceived sender of the argument. The analysis of the resulting corpus of 800 arguments, each annotated by 5 participants, reveals that convincingness is positively correlated with positive emotions (e.g., trust) and negatively correlated with negative emotions (e.g., anger). The appraisal variables disclose the importance of the argument familiarity. For most participants, the content of the argument itself is the primary driver of the emotional response.
Trust Me, I Can Convince You:
The Contextualized Argument Appraisal Framework
Lynn Greschner, Sabine Weber,  and Roman Klinger Fundamentals of Natural Language Processing, University of Bamberg, Germany {lynn.greschner,sabine.weber,roman.klinger}@uni-bamberg.de
1 Introduction
@@
One potential reason is that capturing emotions in text is a challenging task due to their contextualized and implicit nature (Casel et al., 2021; Klinger et al., 2018; Koga et al., 2024; Lee and Lau, 2020). This is particularly true in arguments, which rarely directly express an emotion that is intended to be caused. Figure 1 shows an example in our framework. While the argument appears non-emotional, various factors influence the emotional response: The prior beliefs or personality of the receiver, the relationship with and properties of the sender, or the content of the argument itself. Appraisal theories serve as a suitable basis to model the associated cognitive component (Ellsworth and Scherer, 2003). While they received some attention in general emotion analysis (Troiano et al., 2023), there is neither a psychological framework of argument appraisal, nor is there work in NLP on fine-grained modeling of emotions in arguments.
@@
Next to answering these questions, the main contributions are the novel framework and the associated corpus ContArgA, consisting of 4000 individual annotations for 800 arguments [^1] The data will be made available to the research community under a Creative Commons license..
2 Related Work
@@
2.1 Argument Mining and Convincingness
@@
2.2 Emotion Analysis and Appraisals
@@
Dimension Description Values
Suddenness the argument appears sudden or abrupt to the receiver 1–5
Suppression the receiver tries to shut the argument out of their mind 1–5
Familiarity the argument is familiar to the receiver 1–5
Pleasantness the argument is pleasant for the receiver 1–5
Unpleasantness the argument is unpleasant for the receiver 1–5
Consequencial Importance the argument has important consequences for the receiver 1–5
Positive Consequentiality the argument has positive consequences for the receiver 1–5
Negative Consequentiality the argument has negative consequences for the receiver 1–5
Consequence Manageability the receiver can easily live with the unavoidable consequences of the argument 1–5
Internal Check the consequences of the argument clash with the receiver’s standards and ideals 1–5
External Check the consequences of the argument violate laws or socially accepted norms 1–5
Response urgency the receiver urges to immediately respond to the argument 1–5
Cognitive Effort processing the argument requires a great deal of energy of the receiver 1–5
Argument Internal Check statements in the argument clash with the receiver’s standards and ideals 1–5
Argument External Check statements in the argument violate laws or socially accepted norms 1–5
Table 1: Appraisal dimensions and descriptions used in our study for argument appraisal.
2.3 Emotions in Arguments
@@
3 Contextualized Argument Appraisal Framework
We illustrate the Contextualized Argument Appraisal Framework (CAAF) in Figure 1. CAAF considers arguments within a situational context with the sender, world knowledge, and the self (receiver). The receiver cognitively evaluates, in a synchronized but not necessarily sequential manner, the appraisal, emotion category, and convincingness of the argument.
3.1 Emotions
@@ 3.4 Context
4 Corpus Creation
We now describe our annotation setup based on CAAF, to study the correlations of emotion, appraisal, and convincingness, and pave the way for computational modeling. Screenshots of the study are provided in Appendix E.
4.1 Overview
Figure 2 displays the flowchart of our annotation setup. The study begins with assessing the participant’s emotional state, personality, and stance toward the topic. Subsequently, participants are immersed in a role-playing scenario simulating a town-hall meeting, designed to approximate real-world argumentative contexts and to evoke authentic emotional and cognitive engagement. Exact scenario descriptions and the involvement questions are Appendix D. The role-playing approach, used in NLP to create anntotated corpora Troiano et al. (2024), puts individuals in the shoes of fictional characters.
We ensure emotional involvement in the role-playing scenario, participants describe where they would sit and what they would see in their surroundings. Next, a speaker presents an argument, which participants annotate. They then generate an argument themselves to increase involvement (this data is not evaluated in this paper). Participants further answer demographic and personality trait questions about the speaker. This process is repeated for a second argument within the same role-playing scenario. Lastly, the participants disclose own demographics. Seven attention checks are inlcuded to ensure high data quality, see Appendix E for details.
4.2 Variables
@@
4.2.1 Argument
The following section describes the collected variables on the argument level in detail. All argument variables are annotated with the subjective perception of the argument by the receiver.
@@ 4.2.1 Argument
Our emotion label set comprises Ekman’s basic emotions anger, disgust, fear, joy, sadness, surprise, expanded by shame, guilt, pride, (Greschner and Klinger, 2025), trust and relief (Troiano et al., 2023). Each emotion is rated with its intensity on a 5-point Likert scale. We refer to the emotion with the highest rating as the dominant emotion. In the case of ties participants are requested to chose one. We assess the perceived interest and boredom of the argument on a 5-point Likert scale at a different point in the study and with a different formulation than the emotion questions, see Appendix E for the formulations.
@@
Participants are instructed to provide a short textual explanation of the evoked emotions using a free text field without length restriction.
@@
Participants answer three questions about the influence of the sender, argument content, and receiver on a 5-point Likert scale. The questions are phrased in the format “[The speaker identity, the argument itself, who I am] was important for the emotion I developed”.
@@
We phrase each appraisal dimension as an affirmation, for instance, for response urgency we formulate: “I feel the urge to immediately respond to the argument”. Every appraisal dimension is evaluated on a 5-point Likert scale (not at all, …, extremely).
@@
For each argument, participants provide their perceived degree of manipulative intent. We formulate the question: “Do you feel manipulated by this argument?” and allow answers on a 5-point Likert scale (not at all, …, extremely) [^2] In preliminary studies, we found perceived manipulation to decrease convincingness. In ContArgA, there is a weak, significant negative correlation between manipulation and convincingness, which we leave for future work for a more in-depth analysis..
@@
Similar to the manipulation, participants annotate the perceived convincingness for the question “How convincing is this argument for you?” on a 5-point Likert scale.
4.2.2 Receiver
@@
Participants are asked to provide their stance towards a the topic of the displayed town hall discussion on a 5-point scale (strongly disagree, …, strongly agree). This question is asked once prior to argument exposure and once after each argument, allowing us to measure potential change in the stance after specific arguments. Additionally, participants provide their familiarity with the topic on a 5-point scale (not to be confounded with the appraisal of familiarity with the argument).
@@
Lukin et al. (2017) demonstrate that more conscientious people are rather convinced by emotional arguments. Since we hypothesize that evoked emotions are one of the primary drivers of the argument’s convincingness, we use the Ten-Item Personality Inventory questionnaire (Gosling et al., 2003) to assess the extraversion, agreeableness, conscientiousness, emotional stability, and openness to experiences of the participants.
Argument Topic Emotion–Convincingness Labels
“people simply can’t compete with others in the same professional setting once they reach a certain age, mandatory retirement is necessary for fair competition in a limited job market” Retirement guilt–4; disgust–1; relief–3; sadness–2; surprise–1
“women in polygamous marriages have been shown to be at much greater risk of abuse than those in monogamous marriages” Polygamy disgust–3; sadness–2; anger–3; sadness–3; surprise–2
“there are direct links to drinking bottled water and certain types of disease, even cancer” Plastic bottles fear–4; trust–1; anger–1; guilt–2; surprise–2
Table 2: Examples from the ContArgA corpus from various topics. The emotion & convincingness column contains five individual emotion and convincingness labels for each argument. The labels from different annotators are indicated as emotion–convincingness score.
4.2.3 Sender
In our study setup, the age (old/young) and gender (male/female) of the argument-presenting sender are systematically manipulated. While we are aware that neither age nor gender are binary variables, this simplified setup allows us to investigate whether these properties influence persuasiveness more easily. Each argument is attributed to one single sender, whose age and gender are assigned according to a balanced randomization scheme across topics to ensure approximately equal representation of all four demographic combinations. We present participants with a description of the sender and the argument in the format “A [speaker age] [speaker gender] approaches the microphone and makes a statement: [argument]”.
Following the argument exposure, participants answer the same demographic and TIPI questionnaire questions described in 4.2.2 about their perception of the sender. This approach allows us to examine the interplay between speaker characteristics and audience perception within argumentative discourse. Additionally, the participants can give details about the imagined sender in a free text field. We leave a more in-depth analysis of these variables for future work.
4.3 Crowd-sourcing Details
@@
4.4 Datasets
The UKPConvArgv1 (Habernal and Gurevych, 2016b) and IBM-Rank-30k (Gretz et al., 2019) corpora, hereafter UKP and IBM, serve as sources for the arguments. Both datasets provide isolated arguments and their respective discussion topic. Isolated arguments allow us to investigate the initial argument appraisal without adding noise due to lengthy debates containing multiple arguments. We sample 14 and 26 topics from UKP and IBM, respectively (see Appendix A). For each topic, an equal number of pro and con arguments is sampled. Arguments are manually annotated for comprehensiveness, persuasive attempt, and stance alignment. Importantly, we exclude arguments that point toward the age or gender of the sender, since we manipulate these variables in our study (see Section 4.2.3). See Appendix A for more details on the manual filtering. We exclude 170 annotations due to failed attention checks and randomly select 800 arguments from the 953 valid ones for our study.
5 Data Analysis
Our study results in a corpus of 4,000 individual argument annotations of 800 arguments from 39 topics with 5 individual annotations. Details about the topics can be found in Appendix C.
Table 2 displays example arguments with emotion and convincingness labels. The average length of the arguments is 32.1 tokens, 1.7 sentences. The average value of convincingness is 2.7 (min=1, max=5); the values also depend on the topic, with retirement being most (3.2) and atheism being least convincing (2.0). The frequencies of the dominant emotions are, in descending order: trust (709), surprise (564), sadness (561), anger (521), relief (382), disgust (292), pride (198), joy (193), shame (100), fear (90), guilt (84). The average emotion intensity value across emotions is 1.61, with trust showing the highest (2.02), guilt showing the lowest (1.27) value.
We now turn to the example arguments from ContArgA, displayed in Table 2, in more detail. The first argument evokes 5 different emotions and convincingness labels from 1–4. In contrast, the second argument evokes negative emotions (disgust, sadness, anger) with similar convincingness values, with the exception of surprise.
The second argument discusses the greater risk of abuse of women in polygamous marriages, intuitively evoking negative emotions. Surprise may arise if the receiver is unfamiliar with the topic. For the third argument, we again find negative emotions (fear, anger, guilt). Fear may result from the risk of developing cancer, while a person regularly using plastic bottles could feel guilty. The argument evoking trust is less intuitively clear. These examples illustrate the wide variety of emotions evoked in different participants by the same arguments.
@@ 5 Data Analysis
The contextualized argument appraisal framework is designed to capture individual argument appraisals on divisive topics. We intentionally construct it to elicit differences between receivers of arguments rather than to derive one ‘gold-label’. As expected, the inter-annotator agreement on single variables is therefore low, with a Krippendorff’s Alpha of 0.06 for convincingness and Fleiss’ Kappa of 0.17 for the most dominant emotion category of an argument. These numbers do not suggest a low quality of the data, but justify the need for modeling and analysis under the paradigm of perspectivism.
Figure 3: Emotion Intensities and Convincingness.
Variable r p
Anger -0.222 0.000
Disgust -0.233 0.000
Fear 0.006 1.000
Guilt 0.113 0.000
Joy 0.447 0.000
Pride 0.465 0.000
Relief 0.510 0.000
Sadness -0.125 0.000
Shame -0.052 0.027
Surprise -0.085 0.000
Trust 0.578 0.000
Manipulation -0.104 0.000
Interest 0.691 0.000
Boredom -0.491 0.000
Suddenness -0.344 0.000
Suppression -0.296 0.000
Familiarity 0.336 0.000
Pleasantness 0.576 0.000
Unpleasantness -0.363 0.000
Consequential Importance 0.151 0.000
Positive Consequentiality 0.400 0.000
Negative Consequentiality 0.217 0.000
Consequence Manageability -0.018 1.000
Internal Check -0.105 0.000
External Check -0.366 0.000
Response Urgency -0.249 0.000
Cognitive Effort -0.078 0.000
Argument Internal Check -0.124 0.000
Argument External Check -0.498 0.000
Table 3: Pearson correlations ($r$) and corresponding $p$ values (with Bonferroni correction) for all variables and convincingness.
6 Results
6.1 RQ1: How do discrete emotions and appraisals correlate with argument convincingness?

### v2

@@ Abstract
Emotions that somebody develops based on an argument do not only depend on the argument itself – they are also influenced by a subjective evaluation of the argument’s potential impact on the self. For instance, an argument to ban plastic bottles might cause fear of losing a job for a bottle industry worker, which lowers the convincingness – presumably independent of its content. While binary emotionality of arguments has been studied, such cognitive appraisal models have only been proposed in other subtasks of emotion analysis, but not in the context of arguments and their convincingness. To fill this research gap, we propose the Contextualized Argument Appraisal Framework to model the interplay between the sender, receiver, and argument. We adapt established appraisal models from psychology to argument mining, including argument pleasantness, familiarity, response urgency, and expected effort, as well as convincingness variables. To evaluate the framework and pave the way for computational modeling, we develop a novel role-playing-based annotation setup, mimicking real-world exposure to arguments. Participants disclose their emotion, explain the main cause, the argument appraisal, and the perceived convincingness. To consider the subjective nature of such annotations, we also collect demographic data and personality traits of both the participants and ask them to disclose the same variables for their perception of the argument sender. The analysis of the resulting ContArgA corpus of 4000 annotations reveals that convincingness is positively correlated with positive emotions (e.g., trust) and negatively correlated with negative emotions (e.g., anger). The appraisal variables particularly point to the importance of the annotator’s familiarity with the argument.
Keywords: emotions, appraisals, arguments, convincingness, implicit language
[…]
Trust Me, I Can Convince You: The Contextualized Argument Appraisal Framework and the ContArgA Corpus
Lynn Greschner, Sabine Weber, Roman Klinger
Fundamentals of Natural Language Processing, University of Bamberg, Germany
{firstname.lastname}@uni-bamberg.de
Abstract content
1.  Introduction
@@
One potential reason is that capturing emotions in text is a challenging task due to their contextualized and implicit nature (Casel et al., 2021; Klinger et al., 2018; Koga et al., 2024; Lee and Lau, 2020). This is particularly true in arguments, which rarely directly express an emotion that is intended to be caused. Figure 1 shows an example in our framework. While the argument appears non-emotional, various factors influence the emotional response: prior beliefs of the receiver, their personality, the relationship with the sender and their traits, or the content of the argument itself. Appraisal theories serve as a suitable basis to model the associated cognitive component (Ellsworth and Scherer, 2003). While they received some attention in event-centered emotion analysis (Klinger, 2023), there is neither a psychological framework of argument appraisal, nor is there work in NLP on fine-grained modeling of emotions in arguments.
@@
Next to answering these questions, the main contributions are the novel framework and the associated corpus ContArgA, consisting of 4000 individual annotations for 800 arguments [^1] The data and code for running the annotation study is available here: https://www.uni-bamberg.de/en/nlproc/projects/emcona/..
2.  Related Work
@@
2.1.  Argument Mining and Convincingness
@@
$2.$  Emotion Analysis and Appraisals
@@
Dimension Description
Suddenness the argument appears sudden or abrupt to the receiver
Suppression the receiver tries to shut the argument out of their mind
Familiarity the argument is familiar to the receiver
Pleasantness the argument is pleasant for the receiver
Unpleasantness the argument is unpleasant for the receiver
Consequencial Importance the argument has important consequences for the receiver
Positive Consequentiality the argument has positive consequences for the receiver
Negative Consequentiality the argument has negative consequences for the receiver
Consequence Manageability the receiver can easily live with the unavoidable consequences of the argument
Internal Check the consequences of the argument clash with the receiver’s standards and ideals
External Check the consequences of the argument violate laws or socially accepted norms
Response urgency the receiver urges to immediately respond to the argument
Cognitive Effort processing the argument requires a great deal of energy of the receiver
Argument Internal Check statements in the argument clash with the receiver’s standards and ideals
Argument External Check statements in the argument violate laws or socially accepted norms
Table 1: Appraisal dimensions used to measure argument evaluation in the context of the self.
2.3.  Emotions in Arguments
@@
3.  Contextualized Argument Appraisal Framework
We illustrate the Contextualized Argument Appraisal Framework (CAAF) in Figure 1. CAAF considers arguments within a situational context with the sender and world knowledge, and the self, the receiver. The receiver cognitively evaluates, in a synchronized but not necessarily sequential manner, the appraisal, emotion category, and convincingness of the argument.
3.1.  Emotions
@@ 3.4.  Context
4.  Corpus Creation
We now describe our annotation setup based on CAAF, to study the correlations of emotion, appraisal, and convincingness, and pave the way for computational modeling [^2] Screenshots of the study can be found in Appendix F..
4.1.  Overview
Figure 2 illustrates our annotation setup. The study begins with assessing the participant’s emotional state, personality, and stance toward the topic. Subsequently, participants are immersed in a role-playing scenario simulating a town-hall meeting to approximate a real-world argumentative context. Exact scenario descriptions and the involvement questions are Appendix E.
To ensure that participants do engage with the role-playing, we ask them to describe where they would sit and what they would see in their surroundings. Next, we tell participants that a speaker presents an argument, which they annotate. They then generate an argument themselves to increase involvement (this data is not evaluated in this paper). Participants further answer demographic and personality trait questions about the speaker. This process is repeated for a second argument within the same role-playing scenario. Lastly, the participants disclose own demographics. Seven attention and sanity checks are included in the study, see Appendix F for details.
Annotations
ID Argument Topic Emo. Conv.
1 “people simply can’t compete with others in the same professional setting once they reach a certain age, mandatory retirement is necessary for fair competition in a limited job market” Retirement guilt disgust relief sadness surprise 4 1 3 2 1
2 “women in polygamous marriages have been shown to be at much greater risk of abuse than those in monogamous marriages” Polygamy disgust sadness anger sadness surprise 3 2 3 3 2
3 “there are direct links to drinking bottled water and certain types of disease, even cancer” Plastic bottles fear trust anger guilt surprise 4 1 1 2 2
Table 2: Examples from the ContArgA corpus (argument text source: Habernal and Gurevych, 2016b; Gretz et al., 2019). Emotion and convincingness assessments are from five individual annotators.
4.2.  Variables
@@
4.2.1.  Argument
The following section describes the collected variables on the argument level in detail. All variables are assessed on a 5-point Likert scale if not otherwise stated.
@@ 4.2.1.  Argument
Our emotion label set comprises Ekman’s basic emotions anger, disgust, fear, joy, sadness, surprise, expanded by shame, guilt, pride, (Greschner and Klinger, 2025), trust and relief (Troiano et al., 2023). We refer to the emotion with the highest rating as the dominant emotion. In the case of ties, participants are requested to chose one. We additionally assess the perceived interest and boredom of the argument, see Appendix F for the formulations.
@@
Participants are instructed to provide a short textual explanation of the evoked emotions using a free text field.
@@
Participants answer three questions about the influence of the sender, argument content, and receiver. The questions are phrased in the format “[The speaker identity, the argument itself, who I am] was important for the emotion I developed”.
@@
We phrase each appraisal dimension as an affirmation, e.g., for response urgency we formulate: “I feel the urge to immediately respond to the argument”.
@@
For each argument, participants provide their perceived degree of manipulative intent. We formulate the question: “Do you feel manipulated by this argument?” [^3] In preliminary studies, we found perceived manipulation to decrease convincingness. In ContArgA, there is a weak, significant negative correlation between manipulation and convincingness, which we leave for future work for a more in-depth analysis..
@@
Similar to the manipulation variable, participants annotate the perceived convincingness for the question “How convincing is this argument for you?”
4.2.2.  Receiver
@@
Participants are asked to provide their stance towards a the topic of the displayed town hall discussion. This question is asked once prior to argument exposure and once after each argument, allowing us to measure potential change in the stance after specific arguments. Additionally, participants provide their familiarity with the topic (not to be confounded with the appraisal of familiarity with the argument).
@@
Lukin et al. (2017) demonstrate that more conscientious people are rather convinced by emotional arguments. Since we hypothesize that evoked emotions are one of the primary drivers of the argument’s convincingness, we use the Ten-Item Personality Inventory questionnaire (TIPI, Gosling et al., 2003) to assess the extraversion, agreeableness, conscientiousness, emotional stability, and openness to experiences of the participants.
4.2.3.  Sender
In our study setup, the age (old/young) and gender (male/female) of the argument-presenting sender are systematically manipulated. While we are aware that neither age nor gender are binary variables, this simplified setup allows us to investigate whether these properties influence persuasiveness. Each argument is attributed to one single sender, whose age and gender are assigned according to a balanced randomization scheme across topics to ensure approximately equal representation of all four demographic combinations. We present participants with a description of the sender and the argument in the format “A [speaker age] [speaker gender] approaches the microphone and makes a statement: [argument]”.
Following the argument exposure, participants answer the same demographic and TIPI questionnaire questions described in Section 4.2.2 about their perception of the sender. This approach allows us to examine the interplay between speaker characteristics and audience perception within argumentative discourse. Additionally, the participants can give details about the imagined sender in a free text field. We leave a more in-depth analysis of these variables for future work.
Figure 3: Emotion Intensities and Convincingness.
4.3.  Crowd-sourcing Details
@@
$4.$  Datasets
The UKPConvArgv1 (Habernal and Gurevych, 2016b) and IBM-Rank-30k (Gretz et al., 2019) corpora, hereafter UKP and IBM, serve as sources for the arguments. Both datasets provide isolated arguments and their respective discussion topic. Isolated arguments allow us to investigate the initial argument appraisal without adding noise due to lengthy debates containing multiple arguments. We sample 14 and 26 topics from UKP and IBM, respectively. For each topic, an equal number of pro and con arguments is sampled.
Arguments are manually annotated for comprehensiveness, persuasive attempt, and stance alignment. Importantly, we exclude arguments that point toward the age or gender of the sender, since we manipulate these variables in our study (see Section 4.2.3). We randomly select 800 arguments from the 953 valid ones for our study [^7] Details about the manual filtering in Appendix A..
5.  Data Analysis
Our study results in a corpus of 4,000 individual argument annotations of 800 arguments from 39 topics with 5 individual annotations (170 annotations excluded due to failed attention checks [^8] Details about the topics can be found in Appendix D.).
Table 2 displays example arguments with emotion and convincingness labels. The average length of the arguments is 32.1 tokens, 1.7 sentences. The average value of convincingness is 2.7 (min=1, max=5); the values also depend on the topic, with retirement being most (3.2) and atheism being least convincing (2.0). The frequencies of the dominant emotions are, in descending order: trust (709), surprise (564), sadness (561), anger (521), relief (382), disgust (292), pride (198), joy (193), shame (100), fear (90), guilt (84). The average emotion intensity value across emotions is 1.61, with trust showing the highest (2.02), and guilt showing the lowest (1.27) value.
We now turn to the example arguments from ContArgA, displayed in Table 2, in more detail. The first argument evokes 5 different emotions and convincingness labels from 1–4. In contrast, the second argument evokes negative emotions (disgust, sadness, anger) with similar convincingness values, with the exception of surprise. The second argument discusses the greater risk of abuse of women in polygamous marriages, intuitively evoking negative emotions. Surprise may arise if the receiver is unfamiliar with the topic.
For the third argument, we find negative emotions (fear, anger, guilt). Fear may result from the risk of developing cancer, while a person regularly using plastic bottles could feel guilty. The argument evoking trust is less intuitively clear. These examples illustrate the wide variety of emotions evoked in different participants by the same arguments.
@@ 5.  Data Analysis
The contextualized argument appraisal framework is designed to capture individual argument appraisals on divisive topics. We intentionally construct it to elicit differences between receivers of arguments rather than to derive one ‘gold-label’, leading to 5 individual argument appraisals for each argument. As expected, the inter-annotator agreement on single variables is therefore low, with a Krippendorff’s Alpha of 0.06 for convincingness and Fleiss’ Kappa of 0.17 for the most dominant emotion category of an argument. These numbers do not suggest a low quality of the data. In contrast, they justify the need for modeling and analysis under the paradigm of perspectivism.
Emotions $r$ Appraisals $r$
Trust .58*** Pleas .58***
Relief .51*** Pos Cons .40***
Pride .47*** Fam .34***
Joy .45*** Neg Cons .22***
Guilt .11*** Cons Imp .15***
Fear .01 Cons Man -.02
Shame -.05* Cog Eff -.08***
Surprise -.09*** Int Ch -.11***
Sadness -.13*** Arg Int Ch -.12***
Anger -.22*** Resp Ur -.25***
Disgust -.23*** Sup -.30***
Other Sud -.34***
Interest .69*** Unpleas -.36***
Manipulation -.10*** Ex Ch -.37***
Boredom -.49*** Arg Ex Ch -.50***
Table 3: Pearson correlations ($r$) of emotions and appraisals with the convincingness variable. Significance levels (with Bonferroni correction): * $p<.05$, ** $p<.01$, *** $p<.001$.
6.  Results
In the following, we answer our posed research questions individually.
6.1.  RQ1: How do discrete emotions and appraisals correlate with argument convincingness?